# LLM4ED: Large Language Models for Automatic Equation Discovery

### v1

@@ Abstract
Equation discovery is aimed at directly extracting physical laws from data and has emerged as a pivotal research domain. Previous methods based on symbolic mathematics have achieved substantial advancements, but often require the design of implementation of complex algorithms. In this paper, we introduce a new framework that utilizes natural language-based prompts to guide large language models (LLMs) in automatically mining governing equations from data. Specifically, we first utilize the generation capability of LLMs to generate diverse equations in string form, and then evaluate the generated equations based on observations. In the optimization phase, we propose two alternately iterated strategies to optimize generated equations collaboratively. The first strategy is to take LLMs as a black-box optimizer and achieve equation self-improvement based on historical samples and their performance. The second strategy is to instruct LLMs to perform evolutionary operators for global search. Experiments are extensively conducted on both partial differential equations and ordinary differential equations. Results demonstrate that our framework can discover effective equations to reveal the underlying physical laws under various nonlinear dynamic systems. Further comparisons are made with state-of-the-art models, demonstrating good stability and usability. Our framework substantially lowers the barriers to learning and applying equation discovery techniques, demonstrating the application potential of LLMs in the field of knowledge discovery.
@@ 1 Introduction
Physical laws often follow concise governing equations, which are crucial for our understanding and transformation of the natural world. With the development of artificial intelligence, simulation of the evolution of nonlinear systems through deep learning has gradually emerged [1, 2, 3]. However, these methods are limited by black-box models and lack interpretability. To tackle this issue, equation discovery methods that uncover potential physical laws from observations with explicit mathematic formulas have received increasing attention, which can not only facilitate a deeper understanding of physical processes but also provide domain guidance for data-driven models and enhance their predictive robustness [4, 5]. Moreover, with the governing equation incorporated as physical constraints, neural networks can be equipped with better physical intuition and possess better extrapolation ability [6].
@@
In nonlinear systems, states of interest often follow differential equations, such as ordinary differential equations, in the form of $\dot{\mathbf{x}}=f(\mathbf{x}(t))$, where $\mathbf{x}(t)=\{x_{1}(t),x_{2}(t),...,x_{n}(t)\}^{T}\in\mathbb{R}^{m}$ denotes the state variables with the spatial dimension of $m$. The main objective of equation discovery is to find the explicit expression of $f$. Traditionally, this process was based on first principles, which might require experts in the relevant domain to engage in extensive mathematical derivations. In recent years, data-driven methods are gradually rising because of their superior efficiency and applicability [7, 8]. In particular, SINDy (Sparse Identification of Nonlinear Dynamics) has emerged as an effective method to tackle this challenge [9]. It assumes that the form of $f$ can be simplified as a linear combination of a series of candidate basis functions, where the basis function library is often predetermined based on prior knowledge. With the advantages of high computational efficiency and simple methodology, SINDy has achieved good performance across various fields [10, 11, 12, 13]. Nevertheless, the reliance on prior knowledge inherently constrains the applicability of this approach, rendering it challenging to uncover more intricate representational forms. Concurrently, the progress of numerous intelligent optimization algorithms has contributed to the utilization of symbolic mathematics in identifying governing equations with more flexible forms. EQL (Equation learner) [14, 15] endeavors to utilize the topological structure of networks to represent equations with different combinations and substituting activations with arithmatic operators, such as $+$ and --. An alternative approach seeks to represent equations with expression trees, aiming to discover the optimal equation by optimizing the tree structure. Common optimization methods are based on gradient descent [16, 17, 18], reinforcement learning [19, 20, 21, 22], or evolutionary algorithms [23, 24, 25, 26]. These approaches substantially diminish the reliance on prior physical knowledge, enabling wider application scenarios. However, laborious and intricate algorithm design and coding efforts are required for equation generation and optimization, which is not conducive to wide-scale promotion.
@@
Transformer-based large language models (LLMs) have continuously emerged and have achieved remarkable results in various application domains in recent years [27, 28, 29, 30]. A vast number of trainable parameters and a large diverse training corpus enable LLMs to possess strong generation and reasoning capabilities. Some recent studies have started to explore the potential of LLMs in mathematical reasoning [31], algorithmic optimization [32], and code generation [33], with some even employing LLMs as direct optimizers to tackle black-box optimization challenges [34]. A salient question is whether we can leverage LLMs to automatically complete equation discovery without additional parametric models and optimization processes.
@@
In this paper, we propose a large language model-based framework for automatic equation discovery, as shown in Fig. 1. The equations are first generated with strings format after prompting LLMs with a clear symbol library and problem descriptions. The equations can be seamlessly parsed and transformed into expression trees via the domain tool in symbolic mathematics. During the optimization phase, LLMs can serve as an optimizer to conduct the self-improvement process. LLMs are instructed to conduct local refinements to the historical equations based on the analysis of the inherent relationship between the different combinations of symbols and their performance. On the other hand, well-designed prompts are used to guide LLMs to apply user-defined evolution operators on elite equations, promoting the generation of more diverse equation combinations. These two approaches are iterative and employed in an alternating direction until the optimal equation satisfies the termination conditions. Our framework is testified for uncovering the correct PDE equations in several canonical nonlinear systems and verified that the two optimization approaches of local modification and random evolution have a synergistic effect. In addition, we further validated our framework on sixteen one-dimensional ODE systems, and the results showed that it could achieve comparable performance to the state-of-the-art. Our main contributions are as follows:
@@
  • We validate the efficacy of our framework through a series of experiments on ODEs and PDEs systems. The results demonstrate that its performance is on par with or even better than the state-of-the-art symbolic regression methods, encouraging additional research and application of LLMs in the domain of equation discovery.
@@ 2.1 Symbolic Equation discovery
Symbolic mathematics-based methods can directly uncover the potential relationships between variables from data. With the development of computational equipment and machine learning, these methods have gradually gained increasing attention. Equation discovery tasks typically encompass three phases: generation, evaluation, and optimization. In the generation stage, based on certain context-free grammars [35], equations in mathematical form are transformed into expression trees. The internal nodes of the expression tree are predefined operators (e.g., +,+,-+ , -) and operands (e.g., observations $x$ or constant). By conducting a top-down traversal of the expression, a unique sequential representation can be generated. This representation is more concise and enables more efficient batch generation and gradient-based optimization [19, 36]. Some constraints are carefully designed to generate dimensional consistent expressions and ensure the physical and mathematical rationality. In the evaluation stage, the main focus is to assess the performance of the discovered equations in terms of their fit to the data and complexity. Finally, in the optimization stage, the commonly utilized algorithms mainly include genetic programming [26], gradient descent-based neural network models [14], and recently emerging reinforcement learning models [19, 20]. At the same time, pretrained models based on transformers have gradually emerged [37, 17, 38, 39]. These models are trained on a large amount of data and can directly output the discovered equation results based on the observations, greatly accelerating the inference speed. Evidently, approaches founded on symbolic mathematics necessitate manually designed algorithms in multiple aspects, elevating the learning and application barrier. Conversely, our framework, guided by natural language, significantly streamlines the generation and optimization components, enabling researchers to concentrate solely on the evaluation aspect, where domain expertise is genuinely essential.
@@ 2.2 Large language model for optimization
The powerful language understanding and generation capabilities of large models have led to their extensive application in various fields [27, 28, 29, 30]. Studies have recently demonstrated the feasibility of employing prompt engineering to direct LLMs in addressing optimization problems. One approach is to directly use LLMs as optimizers in a self-improvement updating manner [40, 41]. Taking into account the problem definition and previously generated solutions, LLMs can be directed to refine candidate solutions iteratively. The findings suggest that LLMs possess the capability to progressively improve the generated solutions by building upon the knowledge gained from past optimization results. Other related works attempt to combine LLMs with evolutionary search methods to solve optimizations. Prompts can be designed to instruct LLMs to execute evolutionary algorithms to incrementally enhance the existing solutions within the population. This synergistic combination ultimately leads to the discovery of novel insights and advancements in addressing open research questions, including combinatorial optimization problems like (e.g. traveling salesman problems [34]), multiobjective evolutionary optimization [42], prompt optimization [43], algorithm design [32, 31], game design [44], and evolutionary strategies [45].
Our method pioneered the application of LLMs in the field of equation discovery, constructing a plug-and-play discovery framework. By leveraging natural language, we have seamlessly integrated the self-improvement capabilities of LLMs with evolutionary search techniques, which effectively strikes a balance between exploitation and exploration. The proposed method ensures the stability and efficiency of optimization while finding the optimal equation.
@@ 3.1 Problem overview
The goal of the equation discovery task is to identify an explicit mathematical expression $\mathcal{F}$, defined by mathematical symbols, based on a given set of observations. The true form $\mathbb{F}$ should satisfy
@@
where the state variable $x(t)\in\mathbb{R}^{D}$; $\dot{x}$ refers to the time derivatives; and $\xi$ denotes the possible constants. We aim to find an optimal expression $\mathcal{F}$ that accurately describes the true underlying physical laws in the dynamical system while keeping the form concise. The form of $\mathcal{F}$ may differ slightly for nonlinear systems governed by different types of equations. In this paper, we consider two types of governing equations: PDEs and ODEs. For ODEs, the form of equations can be generated by freely combining symbols from a predefined library, including constants $\xi$. The value of $\xi$ is typically determined using optimization techniques that minimize a specific data fitting metric, such as the mean square error (MSE). For PDEs, the right-hand side of the equation often consists of the combinations of state variables (e.g., $u$) and their spatial derivatives (e.g., $u_{x}$ and $u_{xx}$). Similar to the previous SINDy-based methods [9, 46], we simplify $\mathcal{F}$ to be represented by a linear combination of a series of basis function terms $\Theta(u,x)$. The difference is that function terms can be represented by any combination of symbols without including constants rather than the predefined monomials. Constants only appear as coefficients of the function terms, i.e., $\mathcal{F}\approx\Theta(u,x)\cdot\xi$. The coefficients $\xi$ of the function terms can then be obtained through sparse regression.
In this paper, the expression of $\mathcal{F}$ is generated on the basis of LLMs. We can further empower LLMs with Sympy [47], a domain open-source Python library for symbolic mathematics, to parse the string form and convert it into an expression tree, which facilitates the evaluation of the results.
@@ 3.2 Framework
Our framework employs natural language to guide LLMs in generating and optimizing equations. During the generation process, LLMs draw upon extensive prior training data to produce mathematically reasonable expressions. For optimization, we employ an alternating iterative approach that combines self-improvement and evolutionary search to refine the generated equations. Users are only required to concentrate on establishing appropriate evaluation criteria to precisely evaluate the generated equations. The schematic diagram of our framework is shown in Fig. 2.
@@ 3.3 Initialization
The initial equation population can be generated through LLMs or based on prior knowledge, i.e., manually predefined equations. In this study, we mainly employ prompts to direct LLMs in randomly generating the initial population with a given symbol library. Firstly, LLMs have been trained on extensive text data, enabling them to learn numerous effective equation representations. Consequently, the generated equations generally follow mathematical principles. Secondly, constraints can be established using natural language, thereby preventing the occurrence of equations that violate the specified conditions. For instance, constraints can include restricting the equation length, and frequency of specific symbols, and preventing the generation of invalid nested combinations. Traditionally, implementing these constraints necessitated intricate code, such as generating equations based on probabilistic context-free grammars [48, 49] or subtly modifying probabilities during the symbol sampling process [19].
@@ 3.4 Evaluation
LLMs excels at creative generation based on enormous corpus, but need to be further strengthened with domain tools and human-designed feedback to deal with the symbolc discovery task. Regarding the equation expressions that have been generated in string format, we can employ Sympy [47] to parse and instantiate them as corresponding symbolic expression trees. Before evaluating them, we first need to determine the parameters in the expressions, i.e., constants, and then further score them according to the designed score function.

### v2

# Large Language Models for Automatic Equation Discovery of Nonlinear Dynamics

@@ Abstract
Equation discovery aims to directly extract physical laws from data and has emerged as a pivotal research domain in nonlinear systems. Previous methods based on symbolic mathematics have achieved substantial advancements but often require handcrafted representation rules and complex optimization algorithms. In this paper, we introduce a novel framework that utilizes natural language-based prompts to guide large language models (LLMs) in automatically extracting governing equations from data. Specifically, we first utilize the generation capability of LLMs to generate diverse candidate equations in string form and then evaluate the generated equations based on observations. The best equations are preserved and further refined iteratively using the reasoning capacity of LLMs. We propose two alternately iterated strategies to collaboratively optimize the generated equations. The first strategy uses LLMs as a black-box optimizer to achieve equation self-improvement based on historical samples and their performance. The second strategy instructs LLMs to perform evolutionary operations for a global search. Experiments are conducted on various nonlinear systems described by partial differential equations (PDEs), including Burgers’ equation, the Chafee-Infante equation, and the Navier-Stokes equation. Results demonstrate that our framework can discover correct equations that reveal the underlying physical laws. Further comparisons with state-of-the-art models on extensive ordinary differential equations (ODEs) showcase that the equations discovered by our framework possess physical meaning and better generalization capability on unseen data.
@@ 1 Introduction
Physical laws often follow concise governing equations, which are crucial for our understanding and transformation of the natural world. With the development of artificial intelligence, simulation of the evolution of nonlinear systems through deep learning has gradually emerged [1, 2, 3]. However, these methods are limited by black-box models and lack interpretability. To tackle this issue, equation discovery methods that uncover potential physical laws from observations with explicit mathematical formulas have received increasing attention, which can not only facilitate a deeper understanding of physical processes but also provide domain guidance for data-driven models and enhance their predictive robustness [4, 5]. Moreover, with the governing equation incorporated as physical constraints, neural networks can be equipped with physical intuition and possess better extrapolation ability [6, 7].
@@
In nonlinear systems, states of interest often follow various differential equations, such as ordinary differential equations, in the form of $\dot{\mathbf{x}}=f(\mathbf{x}(t))$, where $\mathbf{x}(t)=\{x_{1}(t),x_{2}(t),...,x_{n}(t)\}^{T}\in\mathbb{R}^{m}$ denotes the state variables with the spatial dimension of $m$. The main objective of equation discovery is to find the explicit expression of $f$. Traditionally, this process was based on first principles, which often require experts in the relevant domain to engage in extensive mathematical derivations. In recent years, data-driven methods are gradually rising because of their superior efficiency and applicability [8, 9]. In particular, SINDy (Sparse Identification of Nonlinear Dynamics) has emerged as an effective method to tackle this challenge [10]. It assumes that the form of $f$ can be simplified as a linear combination of a series of candidate basis functions, where the basis function library is often predetermined based on prior knowledge. With the advantages of high computational efficiency and simple methodology, SINDy has achieved good performance across various fields [11, 12, 13, 14]. Nevertheless, the reliance on prior knowledge inherently constrains the applicability of this approach, rendering it challenging to uncover more intricate representational forms. Concurrently, the progress of numerous intelligent optimization algorithms has contributed to the utilization of symbolic mathematics in identifying governing equations with more flexible forms. EQL (Equation Learner) [15, 16] endeavors to utilize the topological structure of networks to represent equations with different combinations and substitute activations with arithmetic operators, such as $+$ and --. An alternative approach seeks to represent equations with expression trees, aiming to discover the optimal equation by optimizing the tree structure. Common optimization methods are based on gradient descent [17, 18, 19], reinforcement learning [20, 21, 22, 23, 24], or evolutionary algorithms [25, 26, 27]. These approaches substantially diminish the reliance on prior physical knowledge, enabling wider application scenarios. However, laborious and intricate algorithm design and coding efforts are required for equation generation and optimization, which is not conducive to wide-scale promotion.
@@
Transformer-based large language models (LLMs) have continuously emerged and have achieved remarkable results in various application domains in recent years [28, 29, 30]. A vast number of trainable parameters and a large diverse training corpus enable LLMs to possess strong generation and reasoning capabilities. Some recent studies have started to explore the potential of LLMs in mathematical reasoning [31], algorithmic optimization [32], and code generation [33], with some even employing LLMs as direct optimizers to tackle black-box optimization challenges [34]. A salient question is whether we can leverage LLMs to automatically complete equation discovery without additional parametric models and optimization processes.
@@
In this paper, we propose a LLM-based framework for automatic equation discovery, as shown in Fig. 1. Initial equations are first generated in string format after prompting LLMs with a clear symbol library and problem descriptions. The equations can be seamlessly parsed and transformed into expression trees via the domain tool in symbolic mathematics and evaluated based on the score function and data. Elite equations are preserved in the priority queue and incorporated into prompts to guide iterative optimization by LLMs. During the optimization phase, LLMs can serve as an optimizer to conduct the self-improvement process. Some local refinements are applied to the historical equations based on the analysis of the inherent relationship between the combinations of symbols and their performance. In addition, well-designed prompts are used to guide LLMs to apply user-defined evolution operators on elite equations, promoting the generation of more diverse equation combinations. These two approaches are iteratively employed in an alternating manner to refine the structure of generated equations until the optimal equation satisfies the termination conditions. Our framework has been tested for uncovering the correct PDE equations in several canonical nonlinear systems and has verified that the two optimization approaches of local modification and random evolution have a synergistic effect. In addition, we further validated our framework on sixteen one-dimensional ODE systems, and the results showed that it could achieve comparable performance to the state-of-the-art and have better generalization capabilities. Our main contributions are as follows:
@@
  • We validate the efficacy of our framework through a series of experiments on ODEs and PDEs. The results demonstrate that its performance is on par with or even better than the state-of-the-art symbolic regression (SR) methods, especially in generalization capability. The framework encourages more extensive research and application of LLMs in the domain of equation discovery.
@@ 2.1 Symbolic Equation discovery
Symbolic mathematics-based methods can directly uncover the potential relationships between variables from data. With the development of computational equipment and machine learning, these methods have gradually gained increasing attention. Equation discovery tasks typically encompass three phases: generation, evaluation, and optimization. In the generation stage, based on certain context-free s [35, 36], equations in mathematical form are typically transformed into expression trees. The internal nodes of the expression tree are predefined operators (e.g., +,+,-+ , -) and operands (e.g., observations $x$ or constant). By conducting a top-down traversal of the expression, a unique sequential representation can be generated. This representation is more concise and enables more efficient batch generation and gradient-based optimization [20, 37]. Some constraints are carefully designed to generate dimensional consistent expressions and ensure the physical and mathematical rationality. In the evaluation stage, the main focus is to assess the performance of the discovered equations in terms of their fit to the data and complexity. Finally, in the optimization stage, the commonly utilized algorithms mainly include genetic programming [38], gradient descent-based neural network models [15], and recently emerging reinforcement learning models [20, 21]. At the same time, pretrained models based on transformers have gradually emerged [39, 18, 40, 41]. These models are trained on a large amount of data and can directly output the discovered equation results based on the observations, greatly accelerating the inference speed. Evidently, approaches founded on symbolic mathematics necessitate manually designed algorithms in multiple aspects, elevating the learning and application barrier. Conversely, our framework, guided by natural language, significantly streamlines the generation and optimization components, enabling researchers to concentrate solely on the evaluation aspect, where domain expertise is genuinely essential.
@@ 2.2 Large language model for optimization
The powerful language understanding and generation capabilities of large models have led to their extensive application in various fields [28, 29, 30, 42]. Studies have recently demonstrated the feasibility of employing prompt engineering to direct LLMs in addressing optimization problems. One approach is to directly use LLMs as optimizers in a self-improvement updating manner [43, 44]. Taking into account the problem definition and previously generated solutions, LLMs can be directed to refine candidate solutions iteratively. The findings suggest that LLMs possess the capability to progressively improve the generated solutions by building upon the knowledge gained from past optimization results. Other related works attempt to combine LLMs with evolutionary search methods to solve optimization problems. Prompts can be designed to instruct LLMs to execute evolutionary algorithms to incrementally enhance the existing solutions within the population. This synergistic combination ultimately leads to the discovery of novel insights and advancements in addressing open research questions, including combinatorial optimization problems like (e.g. traveling salesman problems [34]), multiobjective evolutionary optimization [45], prompt optimization [46], algorithm design [32, 31], game design [47], and evolutionary strategies [48].
Our method pioneered the application of LLMs in the field of equation discovery, constructing a plug-and-play discovery framework. By leveraging natural language, we have seamlessly integrated the self-improvement capabilities of LLMs with evolutionary search techniques, which effectively strikes a balance between exploitation and exploration. The proposed method ensures the stability and efficiency of optimization while finding the globally optimal equation.
@@ 3.1 Problem overview
The goal of the equation discovery task is to identify an explicit mathematical expression $\mathcal{F}$, defined by mathematical symbols, based on a given set of observations. The true form $\mathcal{F}$ should satisfy
@@
where the state variable $x(t)\in\mathbb{R}^{D}$; $\dot{x}$ refers to the time derivatives; and $\xi$ denotes the possible constants. We aim to find an optimal expression $\mathcal{F}$ that accurately describes the true underlying physical laws in the dynamical system while keeping the form concise. The form of $\mathcal{F}$ may differ slightly for nonlinear systems governed by different types of equations. In this paper, we consider two types of governing equations: PDEs and ODEs. For ODEs, the form of equations can be generated by freely combining symbols from a predefined library, including constants $\xi$. The value of $\xi$ is typically determined using optimization techniques that minimize a specific data fitting metric, such as the mean square error (MSE). For PDEs, the right-hand side of the equation often consists of the combinations of state variables (e.g., $u$) and their spatial derivatives (e.g., $u_{x}$ and $u_{xx}$). For example, the Burgers’ equation is represented as $u_{t}=auu_{x}+bu_{xx}$. Similar to the previous SINDy-based methods [10, 49], we simplify $\mathcal{F}$ to be represented by a linear combination of a series of basis function terms $\Theta(u,x)$. The difference is that function terms can be represented by any combination of symbols without constants rather than the predefined monomials. Constants only appear as coefficients of the function terms, i.e., $\mathcal{F}\approx\Theta(u,x)\cdot\xi$. The coefficients $\xi$ of the function terms can then be obtained through sparse regression.
In this paper, the skeleton of $\mathcal{F}$ is generated and refined by LLMs. We can further empower LLMs with SymPy [50], a domain-specific, open-source Python library for symbolic mathematics, to parse string-form equations and convert them into expression trees, which facilitates the evaluation of data fitting.
@@ 3.2 Framework
Our framework employs natural language to guide LLMs in generating and refining equations, which is shown in Fig. 2. First, LLMs draw upon extensive prior training data, and they tend to produce mathematically reasonable expressions with concrete descriptions of the physical process. Second, we employ an alternating iterative approach that combines self-improvement and evolutionary search to refine the generated equations. Users are only required to concentrate on establishing appropriate evaluation criteria, i.e., the score function, to precisely evaluate the generated equations. Equations with higher scores are used to update a priority queue that retains the top $K$ optimal samples up to the current iteration, while bad samples are discarded. These elite equations are preserved in equation-score format and can be incorporated as in-context examples in prompts to instruct LLMs to generate better-fitting equations. The specific components and procedures of the framework are described below in detail.
@@
3.3 Prompt Engineering
Throughout the entire equation discovery process, the generation and optimization process are instructed with natural language-based prompts, which follow a unified structure in the three processes of initialization, evolution, and self-improvement. The standard format consists of the following components, as shown in Fig. 1.
  • Task descriptions: This part primarily explains the scientific task and defines the symbol library, including operators (e.g., +,+,-+ , - and operands (e.g., $x,const$). If the mechanism of the entire physical process is understood, the physical meanings of different variables can also be provided in this context. This is essential for producing effective and reasonable equations and can significantly reduce the search space.
  • Historical examples: To guide LLMs to generate better-fitting equations, $M$ high-quality equations from past iterations are incorporated as in-context examples in the prompts. The selected equations are chosen from two sources. First, all of the $K$ expressions within the priority queue are used for stable and efficient optimization. Second, we select $M-K$ expressions from the last iteration to maintain sampling diversity. Notably, the presentation of these samples within the prompt varies according to the optimization techniques employed. In the evolution process, only high-quality equations in string format are presented, while in the self-improvement process, historical samples are shown in the form of equation-score pairs.
  • Instructions: This part is primarily aimed at guiding LLMs to generate and refine equations. In the initialization stage, LLMs are required to freely combine symbols from the library to produce equations of arbitrary form and length. In the optimization stage, LLMs are mainly guided to generate refined equations based on historical equations according to different optimization strategies. For the self-improvement process, how to conduct the local modifications should be clearly introduced. The instructions in the LLM-guided evolutionary search should emphasize the concrete implementation details of the selection and evolution of equations.
  • Other hints or constraints: If the relationship of physical variables is available, we can directly describe the requirements for the structure of the generated equations through natural language. In the optimization stage, we can further define local modifications and evolutionary operators and provide possible examples as few-shot prompts. In addition, some hints about the format of outputs can be incorporated here.
The utilized prompts in this paper are demonstrated in Appendix A.
@@ 3.4 Initialization
The initial equation population can be generated through LLMs or based on prior knowledge, i.e., manually predefined equations. In this study, we employ prompts to direct LLMs in randomly generating the initial population with a given symbol library and problem descriptions. First, LLMs have been trained on extensive text data, enabling them to learn numerous effective equation representations. Consequently, the generated equations generally follow mathematical principles. Second, constraints can be established using natural language, thereby preventing the occurrence of equations that violate the specified conditions. For instance, constraints can include restricting the equation length, frequency of specific symbols, and preventing the generation of invalid nested combinations. Traditionally, implementing these constraints necessitated intricate code, such as generating equations based on probabilistic context-free s [35, 51] or subtly modifying probabilities during the symbol sampling process [20].
@@ 3.5 Evaluation
LLMs excels at creative generation based on enormous corpus, but need to be further strengthened with domain tools and human-designed feedback to deal with the symbolic discovery task. Regarding the equation skeletons that have been generated in string format, we can employ Sympy [50] to parse and instantiate them as corresponding symbolic expression trees. Before evaluating them, we first need to determine the parameters in the expressions, i.e., constants, and then further score them according to the designed score function.